OpenStudy (anonymous):

find the rectangular equation for r=4cos theta

OpenStudy (tkhunny):

Multiply by r and see if anything looks familiar.

OpenStudy (anonymous):

got it!

OpenStudy (anonymous):

y^2=-x^2 +4x?

OpenStudy (tkhunny):

Whence came the -x^2?

OpenStudy (perl):

yes thats correct

OpenStudy (anonymous):

sorry i meant x^2+y^2-4x=0

OpenStudy (tkhunny):

Just encouraging you to show your intermediate steps. Now, complete the square in x and tell the REAL story.

OpenStudy (anonymous):

eh its a take home test and the question just asked for the rectangular equation

OpenStudy (anonymous):

thats how they all are

OpenStudy (tkhunny):

Very disappointing. :-)